# Grammatical Error Correction for Low-Resource Languages: The Case of Zarma

### v1

@@ 8 Potential Applications
Our team visited Niamey to present our work to the local Zarma community and gather their feedback. The discussions provided valuable insights into potential applications for our GEC tool and broader language models, highlighting their potential impact on the community.
@@
Feedback from the community highlighted the importance of maintaining accurate written records of folklore, oral histories, and traditional knowledge. The GEC tool can support these efforts by providing a reliable means for transcribing and publishing grammatically accurate texts.
@@ 9 Conclusion and Future Work
In this study, we have demonstrated the potential of leveraging LLMs and MT models to address the critical need for effective grammatical error correction GEC tools in low-resource languages, with a focus on Zarma. The challenges inherent to developing GEC tools for Zarma—such as data scarcity and lack of standardized orthography—necessitated innovative approaches beyond those used in high-resource language contexts. To overcome these challenges, we employed a novel methodology combining three key components: (1) the development of a custom corruption script to generate synthetic training data, thereby mitigating the limited availability of annotated Zarma text; (2) the creation of a human-annotated "Gold Data" set, incorporating expert knowledge of Zarma grammar to provide a valuable benchmark for evaluating model performance on complex errors; and (3) the adaptation and fine-tuning of advanced LLMs and MT models—Gemma 2b and M2M100—for the task of Zarma GEC.
This research establishes a baseline for future work in developing GEC tools for Zarma and other low-resource languages. Future research directions include:
Hybrid Approaches
Exploring hybrid models that combine the precision of rule-based methods with the adaptability of LLMs and the robustness of MT models could yield more effective GEC tools.
Data Augmentation
Further expanding and diversifying the training dataset, possibly through community-driven data collection and annotation efforts, to enhance model performance and generalization capabilities.
Continuous Learning Mechanisms
Incorporating mechanisms for continuous learning and user feedback could enable the models to adapt to evolving language use, idiomatic expressions, and regional dialects.
Resource Optimization
Investigating methods to reduce the computational requirements of advanced models, such as model compression or distillation techniques, to make GEC tools more accessible in resource-constrained environments.
Cross-Lingual Transfer Learning
Applying the insights gained from this study to other low-resource languages, leveraging cross-lingual transfer learning to accelerate the development of GEC tools across diverse linguistic contexts.
@@ Appendix A Further Experiment with other Languages
After obtaining promising results for Zarma using LLM and MT-based approaches, we conducted further experiments to validate the reproducibility of these methods—using the M2M100 and MT5 models. We selected the Bambara language for this experiment because it belongs to a different linguistic family, allowing us to evaluate the performance of the approaches on a language outside the Nilo-Saharan family. We used the Bayelemabaga dataset (Vydrin et al., 2022) for Bambara. The same data preparation process described in the methodology section was followed; however, we excluded any human-annotated data to focus only on spelling correction performance. The results are presented in Table 6.
@@
The Bambara experiment demonstrated that the MT-based approach outperformed the LLMs-based one regarding word-level correction metrics. The MT-based approach achieved a detection rate of 94.64% and a suggestion accuracy of 68.18%. In contrast, the LLMs-based approach detected 87.45% of errors and suggested corrections with 52.91% accuracy. The promising results from the Bambara experiment highlight the potential of both LLMs and MT models to improve GEC for low-resource languages significantly. However, they also emphasize the necessity for continued expanding and diversifying training datasets.
@@ B.2.1 Logical Errors
Logical errors include incorrect verb conjugations, subject-verb agreement issues, incorrect use of grammatical markers, and logical inconsistencies within the sentence. These errors affect the grammatical correctness and logical coherence of the sentence.
@@ B.2.2 Sentence Improvement
Sentence improvement involves enhancing the quality of the sentence by making it more precise, concise, or stylistically appropriate. This category addresses grammatically correct sentences that can be improved for better readability or style.
@@ Appendix C Evaluators Recruitment Process
To ensure a thorough and culturally appropriate manual evaluation of our GEC methods, we recruited five native Zarma speakers to serve as evaluators. The primary criteria for selection was their fluency in Zarma—their ability to read and write proficiently in the language. The evaluators were selected through community outreach in Niamey, Niger, leveraging connections with local educational institutions. All selected evaluators had completed secondary education or higher.
@@
To ensure the reliability of the evaluations, we measured inter-annotator agreement using Fleiss’ kappa coefficient. The resulting kappa value was 0.74, indicating substantial agreement among the evaluators. This suggests that the scoring guidelines were clear and that the evaluators applied them consistently.
@@ Appendix D Training Details
QLoRA GPU Used Lr Loss
Gemma 2b 2 billion Applied NVIDIA P100 $2\times 10^{-4}$ 1.2613
MT5-small 300 million Not Applied NVIDIA T4x2 $2\times 10^{-5}$ 0.0345
M2M100 418 million Not Applied NVIDIA P100 $2\times 10^{-5}$ 0.0214

### v2

@@ 7 Potential Applications
Our team visited Niamey to present our work to the local Zarma community and gather their feedback. The discussions provided valuable insights into potential applications for our GEC tool and broader language models.
@@
Feedback from the community highlighted the importance of maintaining accurate written records of folklore, oral histories, and traditional knowledge. The GEC tool can support these efforts by providing a means for transcribing and publishing grammatically accurate texts.
@@ 8 Conclusion and Future Work
We presented a comparison of rule-based, LLM-based, and MT-based methods for Zarma GEC. The MT-based approach performed best in both controlled and zero-shot scenarios. We confirmed these findings by testing with Bambara language. This research establishes a foundation for future GEC development in other low-resource languages.
Future work includes hybrid models that combine rule-based precision with MT and LLM adaptability, data augmentation to improve performance, continuous learning mechanisms, resource optimization, and cross-lingual transfer learning to support a wider range of languages.
9 Limitations
9.1 Rule-Based Method
Rule-based approaches rely on predefined linguistic rules and dictionaries, which limit their adaptability. They do not handle new or unseen words well. In zero-shot scenarios, where the system encounters words or structures beyond its dictionary, it often fails to correct or detect errors. Additionally, the creation and maintenance of linguistic resources is difficult for low-resource languages, which often lack standardized orthography (Scannell, 2007).
9.2 LLMs Method
LLMs—Gemma 2b and MT5-small—must rely on diverse training data to capture linguistic nuances. Yet, existing LLMs are often trained on data from high-resource languages, which reduces performance in languages like Zarma. A lack of training samples in these languages causes the model to produce lower-quality suggestions (Bender et al., 2021). In addition, fine-tuning LLMs requires significant computational resources, which can be a challenge in low-resource settings.
9.3 MT Method
MT models—M2M100 in our studies—learn from parallel corpora that map incorrect text to correct text. In low-resource languages, such parallel data is often lacking or incomplete (Tiedemann, 2020). Even though multilingual models can leverage patterns from other languages, the absence of sufficient in-language data can lead to contextually imprecise corrections. Moreover, these models require targeted fine-tuning and periodic updates as language use evolves.
@@ Appendix A Further Experiment with other Languages
After obtaining promising results for Zarma using LLM and MT-based approaches, we conducted further experiments to validate the reproducibility of these methods—using the M2M100 and MT5 models. We selected the Bambara language for this experiment because it belongs to a different linguistic family, which allows us to evaluate the performance of the approaches on a language outside the Nilo-Saharan family. We used the Bayelemabaga dataset (Vydrin et al., 2022) for this experiment. The same data preparation process described in the methodology section was followed; however, we excluded any human-annotated data to focus only on spelling correction performance. The results are presented in Table 6.
@@
The Bambara experiment demonstrated that the MT-based approach outperformed the LLMs-based one regarding word-level correction metrics. The MT-based approach achieved a detection rate of 94.64% and a suggestion accuracy of 68.18%. In contrast, the LLMs-based approach detected 87.45% of errors and suggested corrections with 52.91% accuracy. The results from the Bambara experiment highlight the potential of both LLMs and MT models to improve GEC for low-resource languages significantly. However, they also emphasize the necessity for continued expanding and diversifying training datasets.
@@ B.2.1 Logical Errors
Logical errors are inconsistencies in sentence meaning that arise from incorrect verb conjugations, subject-verb agreement, or the misuse of grammatical markers. Unlike grammatical errors, logical errors affect the semantic coherence of a sentence.
@@ B.2.2 Sentence Improvement
Sentence improvement focuses on enhancing readability, precision, and style. While the sentences may already be grammatically correct, improvements ensure better flow and clarity.
@@ Appendix C Evaluators Recruitment Process
To ensure a thorough and culturally appropriate manual evaluation of our GEC methods, we recruited five native Zarma speakers to serve as evaluators. The primary criteria for selection was their fluency in Zarma—their ability to read and write proficiently in the language. The evaluators were selected through community outreach in Niamey, Niger. All selected evaluators had completed secondary education or higher.
@@ Appendix D Training Details
QLoRA GPU Used Lr
Gemma 2b 2 billion Applied P100 $2\times 10^{-4}$
MT5-small 300 million Not Applied T4x2 $2\times 10^{-5}$
M2M100 418 million Not Applied P100 $2\times 10^{-5}$